# Generalised Causal Dantzig

### v1

@@ Abstract
Prediction invariance of causal models under heterogeneous settings has been exploited by a number of recent methods for causal discovery, typically focussing on recovering the causal parents of a target variable of interest. When instrumental variables are not available, the causal Dantzig estimator exploits invariance under the more restrictive case of shift interventions. However, also in this case, one requires observational data from a number of sufficiently different environments, which is rarely available. In this paper, we consider a structural equation model where the target variable is described by a generalised additive model conditional on its parents. Besides having finite moments, no modelling assumptions are made on the conditional distributions of the other variables in the system. Under this setting, we characterise the causal model uniquely by means of two key properties: the Pearson residuals are invariant under the causal model and conditional on the causal parents the causal parameters maximise the population likelihood. These two properties form the basis of a computational strategy for searching the causal model among all possible models. Crucially, for generalised linear models with a known dispersion parameter, such as Poisson and logistic regression, the causal model can be identified from a single data environment.
@@ 1 Introduction
In recent years, causal inference has been used extensively in social sciences (Sobel, 2000), economics (Varian, 2016), and public health (Glass et al., 2013) to determine the effectiveness of interventions and to inform policy decisions. For example, randomised controlled trials are commonly used in clinical studies to determine the causal effect of a treatment on a disease. However, conducting a randomised controlled trial is not always feasible or ethical. So, in some cases, observational studies must be used to infer causal relationships.
One popular method for causal inference in observational studies is propensity score matching (Haukoos and Lewis, 2015), which involves matching individuals who have similar characteristics and differ only in their exposure to a certain treatment or intervention. This technique has been used in various fields, such as education, finance, and healthcare, to determine the causal effect of a particular intervention. The focus of this propensity score matching is on the estimation of the causal effects, rather than the identification of the causal relationships. The latter is also referred to as causal discovery. An alternative approach to causal inference is the use of structural causal models (Pearl, 2009; Bollen and Pearl, 2013). These allow researchers also to model the causal relationships between variables, as well as to estimate the strength of these relationships. Within this framework, a number of methods have been developed for learning the causal graph from observational data, starting from the traditional PC-algorithm (Spirtes et al., 2001) and leading to various alternatives and extensions (Glymour et al., 2019). In many applications, learning the full causal graph is often too ambitious, while the interest is more focussed on discovering the direct causes of a certain variable of interest. This problem, which is also the focus of this paper, is sometimes referred to as causal feature selection (Guyon et al., 2007).
@@
Rothenhäusler et al. (2019) avoids the combinatorial search by limiting the source of heterogeneity. They model perturbations as a system of linear structural equations, where covariates are shifted additively by an instrument. In that scenario, the covariance between the covariates and the residuals under the causal model remains invariant across datasets. Given access to two datasets with sufficiently different shifts, the difference between these invariant covariances provides a moment condition that uniquely identifies the causal model. Estimating the moment condition gives rise to the causal Dantzig estimator. One of the main drawbacks of the causal Dantzig is that the out-of-sample risk under the empirical estimator is often much larger than that of the ordinary least-squares estimator. This is due to the large variance of the causal Dantzig estimator, when the shifts in the different environments are small to moderate.
@@
As a result of causal invariance, causal models enjoy the best out-of-sample prediction guarantees (Arjovsky et al., 2019). Some alternatives to invariant causal prediction focus more closely on this aspect. Rothenhäusler et al. (2021) find that the out-of-sample risk depends on the correlation between the residuals and the instrument. The more uncorrelated they are, the stronger the out-of-sample risk guarantees under unseen covariate shifts, whereby the causal model provides the best guarantees. Assuming access to the instrumental variable that generates the covariate shift, the authors propose to build a regularised estimator, called anchor regression, that regulates the amount of correlation between the residuals and the instrumental variable. While overcoming the issues of causal Dantzig, the main disadvantage of anchor regression is that it requires explicit information about the shift variables. This is often not available in observational studies.
@@
Kania and Wit (2023) take a different view to anchor regression and aim for a solution that lies between the ordinary least-squares solution, which has the best in-sample risk guarantees, and the causal Dantzig solution, which has the best out-of-sample risk guarantees but which can be attained only with enough heterogeneity in the data. They avoid the need of instrumental variables by adding a regularization term to the likelihood that encourages causal invariance. A cross-validation strategy is proposed in order to find the best level of regularization from the available data.
@@
Causal invariance approaches, and their theoretical guarantees, are mostly presented for the case of linear structural equation models, i.e., linear dependences between the variables and Gaussian distributed errors. A small number of studies have considered extensions to more general models. Bühlmann (2020) proposes an extension of anchor regression to the non-linear case, while Kook et al. (2022) consider the case of a transformation model (Hothorn et al., 2014) for the conditional distribution of the target variable given the parents. Similarly, invariant causal prediction has been extended to the non-linear case (Heinze-Deml et al., 2018) and to the case of time-ordered (Pfister et al., 2019) and general response types (Kook et al., 2023). In the latter case, a transformation model is again used for the conditional distribution of the target variable given the parents, while causal invariance is translated into an assumption of invariance of the expected conditional covariance between the environments and the score residuals. As with the linear equation models with Gaussian errors, these extensions are challenged by the requirement of observational data from a number of sufficiently different environments.
@@
This paper proposes a new model-based approach for causal discovery for general response types which does not require multiple environments. In particular, we consider a structural equation model where the target variable is described by a generalised linear model conditional on its parents (McCullagh and Nelder, 1989). The link function can include flexible, possibly non-linear, additive modelling structures. As in Kook et al. (2023), no modelling assumptions are made on the other conditional distributions. Under this setting, we characterise the causal model uniquely by means of two key properties: invariance of the Pearson residuals and maximum of the population likelihood. As the expected squared Pearson residual of the causal model depends only on the dispersion parameter, causal generalised linear models with a known dispersion parameter, such as Poisson or logistic regression, can be identified from a single environment. Here, we will focus on these models. For other models, an estimate of the dispersion parameter can be used.
@@
In section 2, we define the structural equation model formally and the characterization of the causal model describing the target variable. A computationally efficient strategy for detecting the causal model among all possible models from observational data is proposed. We call the resulting model the generalised causal Dantzig estimator of the causal model. Section 3 evaluates the performance of the method on simulation data in the case of Poisson additive regression and performs a comparison with the PC-algorithm. Section 4 shows the applicability of the proposed approach for the identification of the causal determinants of women’s fertility and high earning, using a Poisson and logistic additive regression setting, respectively.
2 Generalised causal Dantzig
@@
In this section, we extend the causal Dantzig estimator. We begin by defining a structural equation model with exponential family target. Then we define a characterization of the causal model, that is unique under certain conditions. Given this characterization, we will define a population and empirical generalised causal Dantzig estimator of the causal model.
@@ 2.1 Structural equation model with exponential family target
    $(\boldsymbol{X},Y)$ is a causal graphical model (Lauritzen, 1996), i.e., it factorises w.r.t. a Directed Acyclic Graph (DAG) that represents the causal relationships between the variables and satisfies a do-calculus under intervention;
@@
    $Y|\boldsymbol{X}_{PA}=\boldsymbol{x}_{PA}\sim EDF\Big{(}b(f_{PA}(\boldsymbol{x}_{PA})),a(\phi)\Big{)}$ with $\boldsymbol{X}_{PA}$ the parents of $Y$ in the causal graph, where
    $EDF(b(\theta),a(\phi))=\exp\bigg{\{}\frac{y\theta-b(\theta)}{a(\phi)}+c(y;\phi)\bigg{\}},\quad\quad\theta=f_{PA}(\boldsymbol{x}_{PA}),$
    and $f_{PA}:\mathbb{R}^{|PA|}\rightarrow\mathbb{R}$ depends on $\boldsymbol{x}\in\mathbb{R}^{p}$ only through $\boldsymbol{x}_{PA}$, while $a(\cdot)$, $b(\cdot)$ and $c(\cdot)$ are known functions that define the distribution in the exponential dispersion family (McCullagh and Nelder, 1989).
Additional regularity conditions on $\boldsymbol{X}$ and $f_{PA}$, such as finite moments for $f_{PA}(\boldsymbol{X})$ in the case of Poisson distributed $Y$, as well as a strictly positive joint density $p(\boldsymbol{X})>0$, should be imposed depending on the the type of exponential family selected. We furthermore assume that the graph is faithful with respect to $Y$.
The distribution of $Y$ conditional on its parents is described by a generalised linear model with a canonical link, while no modelling assumptions are made on the distributions of the predictors $X_{i}$ given their parents. Notable examples in this family, frequently used on a variety of application studies, are Poisson and logistic regression. In both of these cases, $a(\phi)$ is known and equal to 1, while $b(\theta)=e^{\theta}$ for Poisson and $b(\theta)=\log(1+e^{\theta})$ for the logistic regression model. Note that in our definition of the model, linearity is not assumed in the distribution of $Y$ given $\boldsymbol{X}_{PA}$. The special case of $f_{PA}(\boldsymbol{x}_{PA})=\boldsymbol{\beta}_{PA}^{\top}\boldsymbol{x}_{PA}$ for some causal parameters $\boldsymbol{\beta}_{PA}$ may be of interest in practice, but the theoretical derivations do not assume this and apply also to more general models like generalised additive models (Wood, 2017).
2.2 Generalised causal Dantzig characterization of causal model
The structural equation model defined above contains a causal parameter, such as the infinite dimensional functional parameter $f_{PA}$ or the finite dimensional linear parameter $\boldsymbol{\beta}_{PA}$. The identification of this parameter identifies the causal structure governing $Y$. The next theorem describes the properties that characterise the causal model, that is the properties that are satisfied by $f_{PA}$.
@@ 2.2 Generalised causal Dantzig characterization of causal model
Let $(\boldsymbol{X},Y)$ be distributed according to a structural equation model defined in definition 1. Let $\boldsymbol{X}_{PA}$ be the parents of $Y$ in this graph and $f_{PA}$ the true causal model linking $Y$ to its parents. Let $\dot{b}(\theta)$ and $\ddot{b}(\theta)$ be the first and second derivative of the cumulant generator function, respectively.
Then, $f_{PA}$ satisfies the following population conditions:
@@
    $f_{PA}$ is the maximiser of the expected population likelihood of $Y$ and its causal parents,
    $f_{PA}=\arg\max_{f}\mathbb{E}_{\boldsymbol{X}_{PA},Y}\left[\ell_{\boldsymbol{X}_{PA},Y}(f)\right];$
@@
    $f_{PA}$ achieves a perfectly dispersed population Pearson risk,
    $\mathbb{E}_{\boldsymbol{X},Y}\left[\frac{(Y-\dot{b}(f_{PA}(\boldsymbol{X}_{PA})))^{2}}{\ddot{b}(f_{PA}(\boldsymbol{X}_{PA}))}\right]=a(\phi),$
@@
$\mathbb{E}_{\boldsymbol{X}_{PA},Y}\left[\ell_{\boldsymbol{X}_{PA},Y}(f)\right]=\mathbb{E}_{\boldsymbol{X}_{PA},Y}\bigg{[}\frac{Yf(\boldsymbol{X}_{PA})-b(f(\boldsymbol{X}_{PA}))}{a(\phi)}+c(Y;\phi)\bigg{]}$
is maximised by $f_{PA}$.
@@
As for the second property, using the laws of conditional probabilities, we can write the population Pearson risk as
$\mathbb{E}_{\boldsymbol{X},Y}\bigg{[}\frac{(Y-\dot{b}(f_{PA}(\boldsymbol{X}_{PA})))^{2}}{\ddot{b}(f_{PA}(\boldsymbol{X}_{PA}))}\bigg{]}=\mathbb{E}_{\boldsymbol{X}_{PA}}\bigg{[}\mathbb{E}_{Y,\boldsymbol{X}_{{PA}^{C}}}\bigg{[}\frac{(Y-\dot{b}(f_{PA}(\boldsymbol{X}_{PA})))^{2}}{\ddot{b}(f_{PA}(\boldsymbol{X}_{PA}))}\bigg{|}\boldsymbol{X}_{PA}\bigg{]}\bigg{]}$
@@
where ${PA}^{C}$ is the complementary set of the parental set of $Y$, and hence $\boldsymbol{X}_{{PA}^{C}}$ are the non-causal variables. Recalling that the expected values in the population likelihood and Pearson risks are taken w.r.t. $(\boldsymbol{X},Y)$ distributed according to the true causal graphical model, we have in particular that $Y|\boldsymbol{X}_{PA}=\boldsymbol{x}_{PA}\sim EDF\Big{(}b(f_{PA}(\boldsymbol{x}_{PA})),a(\phi)\Big{)}$ for the chosen EDF distribution and $f_{PA}(\boldsymbol{X}_{PA})$ is only a function of $\boldsymbol{X}_{PA}$. Moreover, being a generalised linear model, it holds that $\mathbb{E}[Y|\boldsymbol{X}_{PA}]=\dot{b}(f_{PA}(\boldsymbol{X}_{PA}))$ and $\mathbb{V}ar[Y|\boldsymbol{X}_{PA}]=\ddot{b}(f_{PA}(\boldsymbol{X}_{PA}))a(\phi)$. Using these properties on the inner expectation, we can write the population Pearson risk as
$\mathbb{E}_{\boldsymbol{X},Y}\bigg{[}\frac{(Y-\dot{b}(f_{PA}(\boldsymbol{X}_{PA})))^{2}}{\ddot{b}(f_{PA}(\boldsymbol{X}_{PA}))}\bigg{]}=\mathbb{E}_{\boldsymbol{X}_{PA}}\bigg{[}\frac{\ddot{b}(f_{PA}(\boldsymbol{X}_{PA}))a(\phi)}{\ddot{b}(f_{PA}(\boldsymbol{X}_{PA}))}\bigg{]}=a(\phi).$
@@
The second property in Theorem 2 shows how the Pearson risk is invariant under the causal model, that is it is stable under any distribution of the covariates $\boldsymbol{X}$, including possible changes due to interventions. Unlike the quadratic risk used in the Gaussian case (Kania and Wit, 2023; Rothenhäusler et al., 2019), it is clear how an invariant property for generalised linear models must account also for their inherent heteroscedasticity. This is the role played by the conditional variance at the denominator of the Pearson risk. However, this invariance property can be satisfied by a number of other models. Considering for simplicity a generalised linear model with a linear predictor described by a $p+1$-dimensional vector of parameters $\boldsymbol{\beta}$, the set of solutions for which the Pearson risk condition is satisfied corresponds to a $p$ dimensional manifold in $\mathbb{R}^{p+1}$. The first condition in Theorem 2, stating that the causal model maximises the population likelihood of the marginal model $(\boldsymbol{X}_{PA},Y)$, guarantees an almost sure unique identification of the causal model within this surface, under the conditions described by the next theorem.
@@
Let $S\subset\{1,\ldots,p\}$ indicate a potential set of causal parents of $Y$ in $\boldsymbol{X}$. Let $f_{S}:\mathbb{R}^{|S|}\rightarrow\mathbb{R}$ be the function that depends strictly on $\boldsymbol{X}_{S}$, and that satisfies (i) $f_{S}$ is defined on some functional space $\cal{F}$ with basis $\psi_{S}=\{\psi_{1},\psi_{2},\ldots\}$, i.e., $f_{S}(\boldsymbol{x}_{S})=\boldsymbol{\beta}_{S}^{\top}\psi_{S}(\boldsymbol{x}_{S})$, and (ii) $\boldsymbol{\beta}_{S}$ solves the population likelihood score equations
@@
$S=PA\mbox{ and }f_{S}=f_{PA}\mbox{ a.s.}\quad\mbox{iff}\quad\mathbb{E}_{\boldsymbol{X},Y}\left[\frac{(Y-\dot{b}(\boldsymbol{\beta}_{S}^{\top}\psi_{S}(\boldsymbol{X}_{S})))^{2}}{\ddot{b}(\boldsymbol{\beta}_{S}^{\top}\psi_{S}(\boldsymbol{X}_{S}))}\right]=a(\phi).$
@@
If $S=PA$ and $f_{S}=f_{PA}$ a.s., then Theorem 2 shows how $f_{PA}=\boldsymbol{\beta}_{PA}^{\top}\psi_{PA}$ satisfies the likelihood score equations from the first property and the Pearson risk invariance from the second property. We therefore only need to prove the reverse direction of this theorem.
@@
Except for a null set on the space of distributions on $\boldsymbol{X}_{S}$, under the true graphical model this is true only if $S=PA$ and $f_{S}=f_{PA}$. ∎

### v2

# Causal generalized linear models via Pearson risk invariance

@@ Abstract
Prediction invariance of causal models under heterogeneous settings has been exploited by a number of recent methods for causal discovery, typically focussing on recovering the causal parents of a target variable of interest. Existing methods require observational data from a number of sufficiently different environments, which is rarely available. In this paper, we consider a structural equation model where the target variable is described by a generalized linear model conditional on its parents. Besides having finite moments, no modelling assumptions are made on the conditional distributions of the other variables in the system, and nonlinear effects on the target variable can naturally be accommodated by a generalized additive structure. Under this setting, we characterize the causal model uniquely by means of two key properties: the Pearson risk invariant under the causal model and, conditional on the causal parents, the causal parameters maximize the expected likelihood. These two properties form the basis of a computational strategy for searching the causal model among all possible models. A stepwise greedy search is proposed for systems with a large number of variables. Crucially, for generalized linear models with a known dispersion parameter, such as Poisson and logistic regression, the causal model can be identified from a single data environment. The method is implemented in the R package causalreg.
@@ 1 Introduction
In recent years, causal inference has been used extensively in social sciences (Sobel, 2000), economics (Varian, 2016), and public health (Glass et al., 2013) to determine the effectiveness of interventions and to inform policy decisions. For example, randomized controlled trials are commonly used in clinical studies to determine the causal effect of a treatment on a disease. However, conducting a randomized controlled trial is not always feasible or ethical. So, in some cases, observational studies must be used to infer causal relationships.
One popular method for causal inference in observational studies is propensity score matching (Haukoos and Lewis, 2015), which involves matching individuals who have similar characteristics and differ only in their exposure to a certain treatment or intervention. This technique has been used in various fields, such as education, finance, and healthcare, to determine the causal effect of a particular intervention. The focus of this propensity score matching is on the estimation of the causal effects, rather than the identification of the causal relationships. The latter is also referred to as causal discovery. An alternative approach to causal inference is the use of structural causal models (Pearl, 2009; Bollen and Pearl, 2013). These allow researchers also to model the causal relationships between variables, as well as to estimate the strength of these relationships. Within this framework, a number of methods have been developed for learning the causal graph from observational data, starting from the traditional PC-algorithm (Spirtes et al., 2001) and leading to various alternatives and extensions (Glymour et al., 2019). In many applications, learning the full causal graph is often too ambitious, and, furthermore, interest is more focused on discovering the direct causes of a certain target variable. This problem, which is also the focus of this paper, is sometimes referred to as causal feature selection (Guyon et al., 2007).
@@
Rothenhäusler et al. (2019) avoid the combinatorial search by limiting the source of heterogeneity. They model perturbations as a system of linear structural equations, where covariates are shifted additively by an instrument. In that scenario, the covariance between the covariates and the residuals under the causal model remains invariant across datasets. Given access to two datasets with sufficiently different shifts, the difference between these invariant covariances provides a moment condition that uniquely identifies the causal model. Estimating the moment condition gives rise to the causal Dantzig estimator. One of the main drawbacks of the causal Dantzig is that the out-of-sample risk under the empirical estimator is often much larger than that of the ordinary least-squares estimator. This is due to the large variance of the causal Dantzig estimator, when the shifts in the different environments are small to moderate.
@@
Some alternatives to invariant causal prediction focus more closely on out-of-sample prediction guarantees (Arjovsky et al., 2019). Rothenhäusler et al. (2021) find that the out-of-sample risk depends on the correlation between the residuals and the instrument. The more uncorrelated they are, the stronger the out-of-sample risk guarantees under unseen covariate shifts, whereby the causal model provides the best guarantees. Assuming access to the instrumental variable that generates the covariate shift, the authors propose to build a regularized estimator, called anchor regression, that regulates the amount of correlation between the residuals and the instrumental variable. While overcoming the issues of causal Dantzig, the main disadvantage of anchor regression is that it requires explicit information about the shift variables. This is often not available in observational studies.
@@
Kania and Wit (2023) take a different view to anchor regression and aim for a solution that lies between the ordinary least-squares solution, which has the best in-sample risk guarantees, and the causal Dantzig solution, which has the best out-of-sample risk guarantees but which can be attained only with enough heterogeneity in the data. They avoid the need of instrumental variables by adding a regularization term to the likelihood or pooled risk that encourages causal invariance. A cross-validation strategy is proposed in order to find the best level of regularization from the available data.
@@
Causal invariance approaches, and their theoretical guarantees, are mostly presented for the case of linear structural equation models, i.e., linear dependences between the variables and Gaussian distributed errors. A small number of studies have considered extensions to more general models. Bühlmann (2020) proposes an extension of anchor regression to the nonlinear case, while Kook et al. (2022) consider a transformation model (Hothorn et al., 2014) for the conditional distribution of the target variable given the parents. Similarly, invariant causal prediction has been extended to the nonlinear case (Heinze-Deml et al., 2018) and to the case of time-ordered (Pfister et al., 2019) and general response types (Kook et al., 2024). In the latter case, a transformation model is again used for the conditional distribution of the target variable given the parents, while causal invariance is translated into an assumption of invariance of the expected conditional covariance between the environments and the score residuals. As with the linear equation models with Gaussian errors, these extensions also suffer from the requirement of observational data from a number of sufficiently different environments.
@@
This paper proposes a new model-based approach for causal discovery for general response types which, for some important special cases, does not require multiple environments. In particular, we consider a structural equation model where the target variable is described by a generalized linear model conditional on its parents (McCullagh and Nelder, 1989). The link function can include flexible, possibly nonlinear, additive modelling structures. As in Kook et al. (2024), no modelling assumptions are made on the other conditional distributions. Under this setting, we characterize the causal model uniquely by means of two key properties: invariance of the Pearson risk and achieving the maximum of a conditional population likelihood. As the expected squared Pearson residual of the causal model depends only on the dispersion parameter, causal generalized linear models with a known dispersion parameter, such as Poisson or logistic regression, can be identified from a single environment. Although the results are general, here we will particularly focus on these models.
@@
In section 2, we define the structural equation model formally, including the characterization of the causal model describing the target variable. A statistical approach for detecting the causal model among all possible models from observational data is proposed. A computationally efficient stepwise strategy is presented for systems with many variables. Section 3 evaluates the performance of the method on simulated data in the case of Poisson and logistic regression and performs a comparison with the PC-algorithm. Section 4 considers an illustration of the method on data generated under a controlled experiment, before showing its applicability in realistic settings where the causal model is unknown. In particular, we show how the method can be used to identify the causal determinants of women’s fertility and high earning, respectively.
2 Causal generalized linear model
@@
In this section, we define the proposed method. We begin by defining a structural equation model with exponential family target. Then we define a characterization of the causal model, that is unique under certain conditions. Given this characterization, we will define a population and empirical algorithm for the search of the causal model from observational data.
@@ 2.1 Structural equation model with exponential family target
    $(\boldsymbol{X},Y)$ is a structural causal model (Pearl, 2009);
@@
    $Y|\boldsymbol{X}_{\mbox{\scriptsize PA}}=\boldsymbol{x}_{\mbox{\scriptsize PA}}\sim EDF\Big{(}b(f_{\mbox{\scriptsize PA}}(\boldsymbol{x}_{\mbox{\scriptsize PA}})),a(\phi)\Big{)}$ with $\boldsymbol{X}_{\mbox{\scriptsize PA}}$ the parents of $Y$ in the causal graph, where
    $EDF(b(\theta),a(\phi))=\exp\bigg{\{}\frac{y\theta-b(\theta)}{a(\phi)}+c(y;\phi)\bigg{\}},\quad\quad\theta=f_{\mbox{\scriptsize PA}}(\boldsymbol{x}_{\mbox{\scriptsize PA}}),$
    and $f_{\mbox{\scriptsize PA}}:\mathbb{R}^{|PA|}\rightarrow\mathbb{R}$ depends on $\boldsymbol{x}\in\mathbb{R}^{p}$ only through $\boldsymbol{x}_{\mbox{\scriptsize PA}}$, while $a(\cdot)$, $b(\cdot)$ and $c(\cdot)$ are known functions that define the distribution in the exponential dispersion family (McCullagh and Nelder, 1989).
Additional regularity conditions on $\boldsymbol{X}$ and $f_{\mbox{\scriptsize PA}}$, such as finite moments for $f_{\mbox{\scriptsize PA}}(\boldsymbol{X})$ in the case of Poisson distributed $Y$, as well as a strictly positive joint density $p(\boldsymbol{X})>0$, should be imposed depending on the the type of exponential family selected. We furthermore assume that the graph is faithful with respect to $Y$.
The distribution of $Y$ conditional on its parents is described by a generalized linear model with a canonical link, while no modelling assumptions are made on the distributions of the predictors $X_{i}$ given their parents. Notable examples in this family, frequently used on a variety of application studies, are Poisson and logistic regression. In both of these cases, $a(\phi)$ is known and equal to 1, while $b(\theta)=e^{\theta}$ for Poisson and $b(\theta)=\log(1+e^{\theta})$ for the logistic regression model. In the following, we will consider closely these distributions which belong to the exponential dispersion family and are defined by only one parameter. Besides Poisson and Binomial, other examples include the exponential and logarithmic distributions. Note that in our definition of the model, linearity is not assumed in the distribution of $Y$ given $\boldsymbol{X}_{\mbox{\scriptsize PA}}$. The special case of $f_{\mbox{\scriptsize PA}}(\boldsymbol{x}_{\mbox{\scriptsize PA}})=\boldsymbol{\beta}_{\mbox{\scriptsize PA}}^{\top}\boldsymbol{x}_{\mbox{\scriptsize PA}}$ for some causal parameters $\boldsymbol{\beta}_{\mbox{\scriptsize PA}}$ may be of interest in practice, but the theoretical derivations do not assume this and apply also to more general models like generalized additive models (Wood, 2017).
2.2 Characterization of causal model
The structural equation model defined above contains a causal parameter, such as the infinite dimensional functional parameter $f_{\mbox{\scriptsize PA}}$ or the finite dimensional linear parameter $\boldsymbol{\beta}_{\mbox{\scriptsize PA}}$. The identification of this parameter identifies the causal structure governing $Y$. The next theorem describes the properties that characterise the causal model, that is the properties that are satisfied by $f_{\mbox{\scriptsize PA}}$.
@@ 2.2 Characterization of causal model
Let $(\boldsymbol{X},Y)$ be distributed according to a structural equation model defined in definition 1. Let $\boldsymbol{X}_{\mbox{\scriptsize PA}}$ be the parents of $Y$ in this graph and $f_{\mbox{\scriptsize PA}}$ the true causal model linking $Y$ to its parents. Let $\dot{b}(\theta)$ and $\ddot{b}(\theta)$ be the first and second derivative of the cumulant generator function, respectively.
Then, $f_{\mbox{\scriptsize PA}}$ satisfies the following population conditions:
@@
    $f_{\mbox{\scriptsize PA}}$ is the maximizer of the expected likelihood of $Y$ and its causal parents,
    $f_{\mbox{\scriptsize PA}}=\arg\max_{f}\mathbb{E}_{\boldsymbol{X}_{\mbox{\scriptsize PA}},Y}\left[\ell_{\boldsymbol{X}_{\mbox{\scriptsize PA}},Y}(f)\right];$
@@
    $f_{\mbox{\scriptsize PA}}$ achieves a perfectly dispersed Pearson risk,
    $\mathbb{E}_{\boldsymbol{X},Y}\left[\frac{(Y-\dot{b}(f_{\mbox{\scriptsize PA}}(\boldsymbol{X})))^{2}}{\ddot{b}(f_{\mbox{\scriptsize PA}}(\boldsymbol{X}))}\right]=a(\phi),$
@@
$\mathbb{E}_{\boldsymbol{X}_{\mbox{\scriptsize PA}},Y}\left[\ell_{\boldsymbol{X}_{\mbox{\scriptsize PA}},Y}(f)\right]=\mathbb{E}_{\boldsymbol{X}_{\mbox{\scriptsize PA}},Y}\bigg{[}\frac{Yf(\boldsymbol{X}_{\mbox{\scriptsize PA}})-b(f(\boldsymbol{X}_{\mbox{\scriptsize PA}}))}{a(\phi)}+c(Y;\phi)\bigg{]}$
is maximized by $f_{\mbox{\scriptsize PA}}$.
@@
As for the second property, using the laws of conditional probabilities, we can write the Pearson risk as
$\mathbb{E}_{\boldsymbol{X},Y}\bigg{[}\frac{(Y-\dot{b}(f_{\mbox{\scriptsize PA}}(\boldsymbol{X})))^{2}}{\ddot{b}(f_{\mbox{\scriptsize PA}}(\boldsymbol{X}))}\bigg{]}=\mathbb{E}_{\boldsymbol{X}_{\mbox{\scriptsize PA}}}\bigg{[}\mathbb{E}_{Y,\boldsymbol{X}_{{\mbox{\scriptsize PA}}^{C}}}\bigg{[}\frac{(Y-\dot{b}(f_{\mbox{\scriptsize PA}}(\boldsymbol{X})))^{2}}{\ddot{b}(f_{\mbox{\scriptsize PA}}(\boldsymbol{X}))}\bigg{|}\boldsymbol{X}_{\mbox{\scriptsize PA}}\bigg{]}\bigg{]}$
@@
where ${\mbox{\scriptsize PA}}^{C}$ is the complementary set of the parental set of $Y$, and hence $\boldsymbol{X}_{{\mbox{\scriptsize PA}}^{C}}$ are the non-causal variables. Recalling that the expected values in the likelihood and Pearson risks are taken w.r.t. $(\boldsymbol{X},Y)$ distributed according to the true structural causal model, we have in particular that $Y|\boldsymbol{X}_{\mbox{\scriptsize PA}}=\boldsymbol{x}_{\mbox{\scriptsize PA}}\sim EDF\Big{(}b(f_{\mbox{\scriptsize PA}}(\boldsymbol{x}_{\mbox{\scriptsize PA}})),a(\phi)\Big{)}$ for the chosen EDF distribution and $f_{\mbox{\scriptsize PA}}(\boldsymbol{X})$ is only a function of $\boldsymbol{X}_{\mbox{\scriptsize PA}}$. Moreover, being a generalized linear model, it holds that $\mathbb{E}[Y|\boldsymbol{X}_{\mbox{\scriptsize PA}}]=\dot{b}(f_{\mbox{\scriptsize PA}}(\boldsymbol{X}_{\mbox{\scriptsize PA}}))$ and $\mathbb{V}ar[Y|\boldsymbol{X}_{\mbox{\scriptsize PA}}]=\ddot{b}(f_{\mbox{\scriptsize PA}}(\boldsymbol{X}_{\mbox{\scriptsize PA}}))a(\phi)$. Using these properties on the inner expectation, we can write the Pearson risk as
$\mathbb{E}_{\boldsymbol{X},Y}\bigg{[}\frac{(Y-\dot{b}(f_{\mbox{\scriptsize PA}}(\boldsymbol{X})))^{2}}{\ddot{b}(f_{\mbox{\scriptsize PA}}(\boldsymbol{X}))}\bigg{]}=\mathbb{E}_{\boldsymbol{X}_{\mbox{\scriptsize PA}}}\bigg{[}\frac{\ddot{b}(f_{\mbox{\scriptsize PA}}(\boldsymbol{X}_{\mbox{\scriptsize PA}}))a(\phi)}{\ddot{b}(f_{\mbox{\scriptsize PA}}(\boldsymbol{X}_{\mbox{\scriptsize PA}}))}\bigg{]}=a(\phi).$
@@
The second property in Theorem 2 shows how the Pearson risk is invariant under the causal model, that is it is stable under any distribution of the covariates $\boldsymbol{X}$, including possible changes due to interventions. Unlike the quadratic risk used in the Gaussian case (Kania and Wit, 2023; Rothenhäusler et al., 2019), it is clear how an invariant property for generalized linear models must account also for their inherent heteroscedasticity. This is the role played by the conditional variance at the denominator of the Pearson risk. However, this invariance property can be satisfied by a number of other models. Considering for simplicity a generalized linear model with a linear predictor described by a $p+1$-dimensional vector of parameters $\boldsymbol{\beta}$, the set of solutions for which the Pearson risk condition is satisfied corresponds to a $p$ dimensional manifold in $\mathbb{R}^{p+1}$. The first condition in Theorem 2, stating that the causal model maximizes the expected likelihood of the marginal model $(\boldsymbol{X}_{\mbox{\scriptsize PA}},Y)$, guarantees an almost sure unique identification of the causal model within this surface, under the conditions described by the next theorem.
@@
Let $S\subset\{1,\ldots,p\}$ indicate a potential set of causal parents of $Y$ in $\boldsymbol{X}$. Let $f_{S}:\mathbb{R}^{|S|}\rightarrow\mathbb{R}$ be the function that depends strictly on $\boldsymbol{X}_{S}$, and that satisfies (i) $f_{S}$ is defined on some functional space $\cal{F}$ with basis $\psi_{S}=\{\psi_{1},\psi_{2},\ldots\}$, i.e., $f_{S}(\boldsymbol{x}_{S})=\boldsymbol{\beta}_{S}^{\top}\psi_{S}(\boldsymbol{x}_{S})$, and (ii) $\boldsymbol{\beta}_{S}$ solves the expected likelihood score equations
@@
$S=PA\mbox{ and }f_{S}=f_{\mbox{\scriptsize PA}}\mbox{ a.s.}\quad\mbox{iff}\quad\mathbb{E}_{\boldsymbol{X},Y}\left[\frac{(Y-\dot{b}(\boldsymbol{\beta}_{S}^{\top}\psi_{S}(\boldsymbol{X}_{S})))^{2}}{\ddot{b}(\boldsymbol{\beta}_{S}^{\top}\psi_{S}(\boldsymbol{X}_{S}))}\right]=a(\phi).$
@@
If $S=PA$ and $f_{S}=f_{\mbox{\scriptsize PA}}$ a.s., then Theorem 2 shows how $f_{\mbox{\scriptsize PA}}=\boldsymbol{\beta}_{\mbox{\scriptsize PA}}^{\top}\psi_{\mbox{\scriptsize PA}}$ satisfies the likelihood score equations from the first property and the Pearson risk invariance from the second property. We therefore only need to prove the reverse direction of this theorem.
@@
Except for a null set on the space of distributions on $\boldsymbol{X}_{S}$, under the true structural causal model this is true only if $S=PA$ and $f_{S}=f_{\mbox{\scriptsize PA}}$. ∎